OpenStudy (anonymous):

A ball is launched from a sling shot. After 1 second, the ball is 121 feet in the air. After 2 seconds, it's 224 feet in the air. Using quadratic functions, find the height, in feet, of the ball after 3 seconds in the air.

OpenStudy (amistre64):

you getting any luck with these yet :)

OpenStudy (amistre64):

0a +0b +c = 0 ; c = 0 ---------------------------- a + b = 121 ; a = 121-b ----------------------------- 4a +2b = 224 4(121-b) + 2b = 224 484 -4b +2b = 224 484 - 224 = 2b 260/2 = b ; b = 130 ------------------------------ a + 130 = 121 a = 121 - 130 ; a = -9 ------------------------------------ our quadratic formula becomes: y = -9x^2 +130x y = -9(3)^2+130(3) y = -81 + 390 y = 309 ; the height is 309 ft after 3 seconds.

OpenStudy (amistre64):

typo..... our quadratic EXPRESSION becomes.. lol, my fauxpas